a block of mass 20 kg is kept on a rough inclined plane.angle of repose is 30 degrees.what should be the external force so that it remains at rest when kept on the same inclined plane having angle of inclination of 60 degrees?

Grade:11

1 Answers

Avinash
6 years ago
To keep the block at rest , F min =mg ( sin60° -uk cos 60° ). = 200( 1/2) (√3-u ) =100(√3-u) There fore force required to keep the block at rest is 100( √3-u)